OpenStudy (anonymous):

if a sphere of diameter has 80 units, the diameter divided by 2, how is the surface area of the resulting sphere affected?

OpenStudy (anonymous):

What's the formula for surface area of a sphere? Is it 4pi(r)^2

OpenStudy (anonymous):

i think it is S= 4 pi r^2

OpenStudy (anonymous):

(4pi(20)^2)/4pi(40)^2) = (20 x 20)/(20 x 20 x 2 x 2) = 1/4 So the new surface area is 1/4 of the original one

OpenStudy (anonymous):

thank you